# Finite energy well-posedness for nonlinear Schrödinger equations with non-vanishing conditions at infinity

### v3

@@ 1.4. The energy-critical equation
The well-posedness of (1.17) is not addressed by Theorem 1.3. Local well-posedness for small data is introduced in [22, Theorem 1.3]. Furthermore, note that the cubic-quintic equation
@@
with $\alpha_{1},\alpha_{3},\alpha_{5}>0$, $\alpha_{3}^{2}-4\alpha_{1}\alpha_{5}>0$ and far-field (1.2) is known to be globally well-posed in the respective energy space due to [44]. The cubic-quintic nonlinearity considered satisfies Assumption 1.5 and is such that $F(1)=0$ and $F(\rho)>0$ for all $\rho>1$. The authors rely on the affine structure of the respective energy space for $d=3$, the perturbative approach introduced in [63, 64] and the well-posedness of the energy-critical nonlinear Schrödinger equation with trivial far-field [18]. This approach can be adapted to show global well-posedness of (1.17). More precisely, it is straightforward to update the perturbative argument, see [44, Eq. (1.14) and (1.15)] to the respective problem for (1.17), see also (4.3).
@@ 1.6. Notations
We fix some notations. We denote by $\mathcal{L}^{d}$ the $d$-dimensional Lebesgue measure. The usual Lebesgue spaces are denoted by $L^{p}(\Omega)$ for $\Omega\subset{\mathbf{R}}^{d}$ and Lebesgue exponent $p\in[1,\infty]$. Sobolev spaces are denoted by $H^{s}({\mathbf{R}}^{d})$ with norm $\|f\|_{H^{s}({\mathbf{R}}^{d})}=\|\left\langle\xi\right\rangle^{s}\hat{f}\|_{L^{2}}$, where $\hat{f}$ denotes the Fourier transform. For $k\in{\mathbf{Z}}$ and $r\in[1,\infty]$, we denote $W^{k,r}$ for the Sobolev space with norm $\|f\|_{W^{k,r}}=\sum_{|\alpha|\leq k}\|D^{\alpha}f\|_{L^{r}({\mathbf{R}}^{d})}$. Mixed space-time Lebesgue or Sobolev spaces are indicated by $L^{p}(I;W^{k,r}({\mathbf{R}}^{d}))$. To shorten notations, we write $L_{t}^{p}W_{x}^{k,r}$ when there is no ambiguity. Further, $C(I;H^{s}({\mathbf{R}}^{d}))$ and $C(I;\mathbb{E}({\mathbf{R}}^{d}))$ denote the space of continuous $H^{s}$- and $\mathbb{E}$-valued functions respectively. Finally, $C>0$ denotes any absolute constant.
@@ 2. The energy space and the linear propagator
In the present paper, we define the energy space $\mathbb{E}$ as in (1.9), see also [17, Section 2]. For the GP equation (1.6), being the prototype for (1.1) with non-vanishing far-field, the energy space considered in [23, 24] consists of the set of wave-functions of finite Ginzburg-Landau energy $\mathcal{E}_{\mathrm{GL}}(\psi)$ is more convenient when dealing with general nonlinearities $f$. In general, $\mathbb{E}\subset\{\mathcal{H}(\psi)<+\infty\}$ while the converse inclusion only holds under further assumptions on $f$. The energy space $(\mathbb{E},d_{\mathbb{E}})$, endowed with the metric (1.12) can be shown to be a complete metric space and be thought of as the analogue of $H^{1}$ for NLS equations with trivial far-field. However, $\mathbb{E}$ is not a vector space and wave functions $\psi\in\mathbb{E}({\mathbf{R}}^{d})$ may exhibit oscillations at spatial infinity, in particular for low dimensions. A suitable characterisation of the energy space and the action of the Schrödinger semigroup on $\mathbb{E}$ is essential for the subsequent well-posedness theory. Despite many of the facts proven here can be found in the literature [23, 24, 17], we provide a self-contained characterisation of the energy space $\mathbb{E}$.
We start by proving that any $\psi\in\mathbb{E}({\mathbf{R}}^{d})$ can be decomposed as sum of a $X^{1}$-function and an $H^{1}$-function, where the Zhidkov space $X^{1}({\mathbf{R}}^{d})$ is defined in (1.11). Following [23, Lemma 1], let $\chi\in C_{c}^{\infty}({\mathbf{C}},{\mathbf{R}})$ be a smooth cut-off function such that
(2.1) $\chi(z)=1\quad|z|\leq 2,\qquad\chi(z)\leq 1\quad z\in{\mathbf{C}},\qquad\operatorname{supp}(\chi)\subset B_{3}(0).$
@@
(2.2) $\psi_{\infty}:=\chi(\psi)\psi,\qquad\psi_{q}:=(1-\chi(\psi))\psi$
for which we have the following bounds.
@@
$\|\psi_{\infty}\|_{X^{1}({\mathbf{R}}^{d})}\leq C\left(1+\sqrt{\mathcal{E}(\psi)}\right),\quad\|\psi_{q}\|_{H^{1}({\mathbf{R}}^{d})}\leq C\sqrt{\mathcal{E}(\psi)}.$
@@
Given the decomposition (2.2), we show that $\psi_{\infty}\in X^{1}({\mathbf{R}}^{d})$. As $\psi_{\infty}\in L^{\infty}({\mathbf{R}}^{d})$ it suffices to check that
$\|\nabla\psi_{\infty}\|_{L^{2}({\mathbf{R}}^{d})}=\|\chi(\psi)\nabla\psi+\psi\chi^{\prime}(\psi)\nabla\psi\|_{L^{2}({\mathbf{R}}^{d})}\leq C\|\nabla\psi\|_{L^{2}({\mathbf{R}}^{d})}.$
The bound $\psi_{q}\in L^{2}({\mathbf{R}}^{d})$ follows from the pointwise inequality $|\psi_{q}|\leq C\left||\psi_{q}|-1\right|$ valid on the support of $1-\chi(\psi)$ and
$\|\nabla\psi_{q}\|_{L^{2}({\mathbf{R}}^{d})}\leq C\|\nabla\psi\|_{L^{2}({\mathbf{R}}^{d})}.$
@@
by means of Minkowski’s inequality. It remains to prove that $(\mathbb{E},d_{\mathbb{E}})$ is a complete metric space. One readily verifies that $d_{\mathbb{E}}$ defines a distance function on $\mathbb{E}({\mathbf{R}}^{d})$. To check that $(\mathbb{E},d_{\mathbb{E}})$ is complete, let $\{\psi_{n}\}_{n}\subset\mathbb{E}$ be a Cauchy sequence w.r.t to $d_{\mathbb{E}}$. Then, there exists $\psi\in X^{1}+H^{1}$ such that $\psi_{n}\rightarrow\psi$ strongly in $X^{1}+H^{1}$. By lower semi-continuity of norms and (1.10) it follows that $\psi\in\mathbb{E}$. ∎
@@ 2.1. The structure of the energy space depending on the dimension
The structure of the energy space $\mathbb{E}({\mathbf{R}}^{d})$ is sensitive to the dimension $d$. To illustrate this, we recall the following fact. Let $\phi\in\mathcal{D}^{\prime}({\mathbf{R}}^{d})$, if $\nabla\phi\in L^{p}({\mathbf{R}}^{d})$ for some $p<d$, then there exists $c\in{\mathbf{C}}$ such that $\phi-c\in L^{p^{\ast}}({\mathbf{R}}^{d})$, where $p^{\ast}=\frac{dp}{d-p}$, see for instance [38, Theorem 4.5.9]. Hence, if $\psi\in\mathbb{E}({\mathbf{R}}^{3})$, then $\psi$ admits a decomposition $\psi=c+v$ where $c\in{\mathbf{C}}$ with $|c|=1$ and $v\in\dot{H}^{1}({\mathbf{R}}^{3})$, where
@@
denotes the completion of $C_{0}^{\infty}({\mathbf{R}}^{3})$ with respect to the $L^{2}$ norm of the gradient. This observation allows for a equivalent definition of $\mathbb{E}({\mathbf{R}}^{3})$. As in [23, Section 4], we introduce
(2.5) $\mathcal{F}_{c}=\left\{v\in\dot{H}^{1}({\mathbf{R}}^{3})\,:\,|v|^{2}+2\operatorname{Re}(c^{-1}v)\in L^{2}({\mathbf{R}}^{3})\right\}.$
One readily checks that
$\tilde{\delta}(u,v)=\|\nabla u-\nabla v\|_{L^{2}({\mathbf{R}}^{3})}+\||u|^{2}+2\operatorname{Re}(c^{-1}u)-2\operatorname{Re}(c^{-1}v)-|v|^{2}\|_{L^{2}({\mathbf{R}}^{3})}$
defines a distance function on $\mathcal{F}_{c}$. One has the following characterisation given by [23, Proposition 4.1].
@@
In [23], the Proposition is stated for $(\mathbb{E}_{\mathrm{GL}},d_{\mathbb{E}_{\mathrm{GL}}})$. We prove below, see Lemma 2.6, that the two metric spaces can be identified and the equivalence of the metrics.
@@
We observe that the connected components of $\mathbb{E}({\mathbf{R}}^{3})$ are given by $c+\mathcal{F}_{c}({\mathbf{R}}^{3})$ for $c\in{\mathbf{C}}$ with $|c|=1$. The energy space $\mathbb{E}({\mathbf{R}}^{3})$ is an affine space and the far-field behavior is determined by $c$ corresponding to a phase shift. The affine structure of the energy space allows for an alternative approach to solve the Cauchy Problem for $d=3$, as observed in [23, Remark 4.5] for (1.6) and exploited in [44] for cubic-quintic nonlinearities and far-field behavior (1.2).
@@
The $2D$ energy space $\mathbb{E}({\mathbf{R}}^{2})$ lacks an affine structure due to non-trivial oscillations at spatial infinity. Indeed, unbounded phase oscillations at spatial infinity may occur, e.g. $\psi(x)=\mathrm{e}^{\mathrm{i}(2+\log|x|)^{\beta}}$ with $\beta<\frac{1}{2}$ is such that $\psi\in\mathbb{E}({\mathbf{R}}^{2})$, see [23, Remark 4.2]. Moreover, the metric space $(\mathbb{E}({\mathbf{R}}^{2}),d_{\mathbb{E}})$ is not separable. We refer to Remark 2.7 for a detailed discussion and a weakened topology for which $\mathbb{E}({\mathbf{R}}^{2})$ is connected and separable.
@@ 2.2. The Hamiltonian for wave-functions in the energy space
We observe that if $\psi\in\mathbb{E}({\mathbf{R}}^{d})$, then it follows from the Chebychev inequality that
(2.8) $\mathcal{L}^{d}(\{\left||\psi|-1\right|>\delta\}\leq\frac{1}{\delta^{2}}\||\psi|-1\|_{L^{2}({\mathbf{R}}^{d})}^{2},$
where $\mathcal{L}^{d}$ denotes the $d$-dimensional Lebesgue measure. Consequently, if $\eta\in C_{c}^{\infty}([0,\infty))$ with $\operatorname{supp}(\eta)\subset[\frac{1}{2},\frac{3}{2}]$ such that
then for all $\psi\in\mathbb{E}({\mathbf{R}}^{d})$ the support of $(1-\eta(|\psi|))$ is of finite Lebesgue measure
(2.10) $\mathcal{L}^{d}(\operatorname{supp}(1-\eta(|\psi|)))\leq\frac{1}{4}\mathcal{E}(\psi).$
The following inequality turns out to be handy for applications in the sequel. For any $q\in[1,\infty)$ there exists $C_{q}>0$ such that for all $\phi\in L_{\operatorname{loc}}^{1}({\mathbf{R}}^{2})$ with $\mathcal{L}^{2}(\operatorname{supp}(\phi))<+\infty$ and $\nabla\phi\in L^{2}({\mathbf{R}}^{2})$ it holds
@@
see for instance [17, Proof of Lemma 2.1]. For $\psi\in\mathbb{E}({\mathbf{R}}^{2})$, applying (2.11) to $\phi=\psi(1-\eta(|\psi|))$ yields $\psi(1-\eta(|\psi|))\in L^{q}({\mathbf{R}}^{2})$ for any $q\in[1,\infty)$. Indeed, it suffices to check that
$\nabla\left(\psi(1-\eta(|\psi|))\right)=(1-\eta(\psi))\nabla\psi-\eta^{\prime}(\psi)\psi\nabla|\psi|\in L^{2}({\mathbf{R}}^{2})$
since $(1-\eta(\psi))\in L^{\infty}({\mathbf{R}}^{2})$, $\psi\eta^{\prime}(\psi)\in L^{\infty}({\mathbf{R}}^{2})$ as well as $|\nabla|\psi||\leq|\nabla\psi|$ a.e. on ${\mathbf{R}}^{2}$.
Under Assumption 1.1, the functional $\mathcal{H}(\psi)$, introduced in (1.10), is bounded for all $\psi\in\mathbb{E}({\mathbf{R}}^{d})$.
@@
In view of (K1) Assumption 1.1, it suffices to use a Taylor expansion of $F$ in a small neighborhood $\mathcal{O}$ of $1$ to show that there exist $C,C^{\prime}>0$ such that
@@
for all $x\in{\mathbf{R}}^{d}$ such that $|\psi|^{2}\in\mathcal{O}$. Let $\delta>0$ such that $B(1,\delta)\subset\mathcal{O}$ and $\eta_{\delta}(r):=\eta(\frac{r}{\delta})$ with $\eta$ as in (2.9) and $\psi\in\mathbb{E}({\mathbf{R}}^{d})$, then
@@
where we used (K2) Assumption 1.1 in the last inequality. To control the second term, we consider separately the cases $d=2,3$. For $d=3$, Proposition 2.2 yields that there exists $c\in{\mathbf{C}}$ with $|c|=1$ and $v\in\mathcal{F}_{c}({\mathbf{R}}^{3})$ such that $\psi=c+v$ and
@@
where we used (2.10) in the second last inequality and that $20<\alpha<$ 2 for $d=3$. For $d=2$, one has that
@@
Next, we identify suitable conditions on $f$ under which the converse inclusion, namely $\{\psi:|\mathcal{H}(\psi)|<+\infty\}\subset\mathbb{E}({\mathbf{R}}^{d})$, holds true. First, we treat the particular case of the Gross-Pitaevskii equation (1.6) for which $\mathcal{H}(\psi)=\mathcal{E}_{\mathrm{GL}}(\psi)$ and thus $\mathbb{E}_{\mathrm{GL}}({\mathbf{R}}^{d})=\{\mathcal{H}(\psi)<+\infty\}$, see (1.7) and (1.8) respectively. It has been shown in [23], see also [24], that $(\mathbb{E}_{\mathrm{GL}},d_{\mathbb{E}_{\mathrm{GL}}})$ with
@@
is a complete metric space. It is pointed out in [17, p.13] without proof that $\mathbb{E}=\mathbb{E}_{\mathrm{GL}}$ with equivalence of the respective metrics. We provide a proof for the sake of completeness.
@@
Lemma 2.6 allows to infer the topological properties of $(\mathbb{E},d_{\mathbb{E}})$ from the results for $(\mathbb{E}_{\mathrm{GL}}({\mathbf{R}}^{d}),d_{\mathbb{E}_{\mathrm{GL}}})$ in [23, 24]. For instance, the functional $\mathcal{E}$ measures the distance to the circle of constants $S^{1}=\{\psi\in\mathbb{E}:\mathcal{E}(\psi)=0\}$ for $d=3$ but not for $d=2$. Indeed, it follows from Lemma 2.6 and [23, Proposition 4.3] that there exists $A>0$ such that for every $\psi\in\mathbb{E}({\mathbf{R}}^{3})$,
@@
If $d=2$, there exists a sequence $\{\psi_{n}\}$ in $\mathbb{E}({\mathbf{R}}^{2})$ such that $\mathcal{E}(\psi_{n})\rightarrow 0$ but $d_{\mathbb{E}}(\psi_{n},S^{1})\geq c_{0}>0$. Note that the complete metric space $(\mathbb{E}_{\mathrm{GL}}({\mathbf{R}}^{2}),d_{\mathbb{E}_{\mathrm{GL}}})$ lacks an affine structure and to be separable. In [24] a detailed characterisation of $\mathbb{E}_{\mathrm{GL}}({\mathbf{R}}^{d})$ including a manifold structure for $\mathbb{E}_{\mathrm{GL}}({\mathbf{R}}^{d})$ is provided. The connected components are characterised by [24, Theorem 1.8] and [24, Proposition 1.10]. A (strictly) weaker topology [24, p. 140] induced by the metric

### v4

@@ 1.4. The energy-critical equation
The well-posedness of (1.17) is not addressed by Theorem 1.3. Local well-posedness for small data is introduced in [25, Theorem 1.3]. Furthermore, note that the cubic-quintic equation
@@
with $\alpha_{1},\alpha_{3},\alpha_{5}>0$, $\alpha_{3}^{2}-4\alpha_{1}\alpha_{5}>0$ and far-field (1.2) is known to be globally well-posed in the respective energy space due to [47]. The cubic-quintic nonlinearity considered satisfies Assumption 1.5 and is such that $F(1)=0$ and $F(\rho)>0$ for all $\rho>1$. The authors rely on the affine structure of the respective energy space for $d=3$, the perturbative approach introduced in [67, 68] and the well-posedness of the energy-critical nonlinear Schrödinger equation with trivial far-field [19]. This approach can be adapted to show global well-posedness of (1.17). More precisely, it is straightforward to update the perturbative argument, see [47, Eq. (1.14) and (1.15)] to the respective problem for (1.17), see also (4.3).
@@ 1.6. Notations
We fix some notations. We denote by $\mathcal{L}^{d}$ the $d$-dimensional Lebesgue measure. The usual Lebesgue spaces are denoted by $L^{p}(\Omega)$ for $\Omega\subset{\mathbf{R}}^{d}$ and Lebesgue exponent $p\in[1,\infty]$. Sobolev spaces are denoted by $H^{s}({\mathbf{R}}^{d})$ with norm $\|f\|_{H^{s}({\mathbf{R}}^{d})}=\|\left\langle\xi\right\rangle^{s}\hat{f}\|_{L^{2}}$, where $\hat{f}$ denotes the Fourier transform. For $k\in{\mathbf{Z}}$ and $r\in[1,\infty]$, we write $W^{k,r}$ for the Sobolev space with norm $\|f\|_{W^{k,r}}=\sum_{|\alpha|\leq k}\|D^{\alpha}f\|_{L^{r}({\mathbf{R}}^{d})}$. Mixed space-time Lebesgue or Sobolev spaces are indicated by $L^{p}(I;W^{k,r}({\mathbf{R}}^{d}))$. To shorten notations, we write $L_{t}^{p}W_{x}^{k,r}$ when there is no ambiguity. Further, $C(I;H^{s}({\mathbf{R}}^{d}))$ and $C(I;\mathbb{E}({\mathbf{R}}^{d}))$ denote the space of continuous $H^{s}$- and $\mathbb{E}$-valued functions respectively. Finally, $C>0$ denotes any absolute constant.
@@ 2. The energy space and the linear propagator
In the present paper, we define the energy space $\mathbb{E}$ as in (1.9), see also [17, Section 2]. For the GP equation (1.6), being the prototype for (1.1) with non-vanishing far-field, the energy space considered in [26, 27] consists of the set of wave-functions of finite Ginzburg-Landau energy $\mathcal{E}_{\mathrm{GL}}(\psi)$ is more convenient when dealing with general nonlinearities $f$. In general, $\mathbb{E}\subset\{\mathcal{H}(\psi)<+\infty\}$ while the converse inclusion only holds under further assumptions on $f$. The energy space $(\mathbb{E},d_{\mathbb{E}})$, endowed with the metric (1.12) can be shown to be a complete metric space and may be regarded as the analogue of $H^{1}$ for NLS equations with trivial far-field. However, $\mathbb{E}$ is not a vector space and wave functions $\psi\in\mathbb{E}({\mathbf{R}}^{d})$ may exhibit oscillations at spatial infinity, in particular for low dimensions. A suitable characterization of the energy space and the action of the Schrödinger semigroup on $\mathbb{E}$ is essential for the subsequent well-posedness theory. Although many of the facts proven here can be found in the literature [26, 27, 17], we provide a self-contained characterization of the energy space $\mathbb{E}$.
To that end, we rely on decompositions of functions by means of smooth cut-off functions that we introduce. Let $\chi\in C_{c}^{\infty}([0,\infty))$ be a smooth cut-off function such that
(2.1) $\mathbf{1}_{[0,2]}(r)\leq\chi(r)\leq\mathbf{1}_{[0,3)}(r).$
@@
(2.2) $\psi_{\mathrm{bd}}:=\chi(|\psi|)\psi,\qquad\psi_{\mathrm{int}}:=(1-\chi(|\psi|))\psi,$
Further, let $\eta\in C_{c}^{\infty}([0,\infty))$ with $\operatorname{supp}(\eta)\subset[\frac{1}{2},\frac{3}{2}]$ such that
We observe that if $\psi\in\mathbb{E}({\mathbf{R}}^{d})$, then it follows for all $c>0$ from the Chebychev inequality that, for all $c>0$,
(2.4) $\mathcal{L}^{d}(\{\left||\psi|-1\right|>c\}\leq\frac{1}{c^{2}}\||\psi|-1\|_{L^{2}({\mathbf{R}}^{d})}^{2},$
where $\mathcal{L}^{d}$ denotes the $d$-dimensional Lebesgue measure. Consequently, for all $\psi\in\mathbb{E}({\mathbf{R}}^{d})$ the supports of $(1-\chi(|\psi|))$ and $(1-\eta(|\psi|))$ are of finite Lebesgue measure
(2.5) $\mathcal{L}^{d}(\operatorname{supp}(1-\chi(|\psi|)))\leq\mathcal{E}(\psi),\qquad\mathcal{L}^{d}(\operatorname{supp}(1-\eta(|\psi|)))\leq 16\mathcal{E}(\psi).$
Following [26, Lemma 1], we start by proving that any $\psi\in\mathbb{E}({\mathbf{R}}^{d})$ can be decomposed as sum of a $X^{1}$-function and an $H^{1}$-function, where the Zhidkov space $X^{1}({\mathbf{R}}^{d})$ is defined in (1.11).
@@
$\|\psi_{\mathrm{bd}}\|_{X^{1}({\mathbf{R}}^{d})}\leq C\left(1+\sqrt{\mathcal{E}(\psi)}\right),\quad\text{and}\quad\|\psi_{\mathrm{int}}\|_{H^{1}({\mathbf{R}}^{d})}\leq C\sqrt{\mathcal{E}(\psi)}.$
@@
Given the decomposition (2.2), we show that $\psi_{\mathrm{bd}}\in X^{1}({\mathbf{R}}^{d})$. As $\psi_{\mathrm{bd}}\in L^{\infty}({\mathbf{R}}^{d})$ it suffices to check that
$\|\nabla\psi_{\mathrm{bd}}\|_{L^{2}({\mathbf{R}}^{d})}=\|\chi(|\psi|)\nabla\psi+\psi\chi^{\prime}(|\psi|)\nabla|\psi|\|_{L^{2}({\mathbf{R}}^{d})}\leq C\|\nabla\psi\|_{L^{2}({\mathbf{R}}^{d})},$
where we used $|\nabla|\psi||\leq|\nabla\psi|$ a.e. on ${\mathbf{R}}^{d}$. The bound $\psi_{\mathrm{int}}\in L^{2}({\mathbf{R}}^{d})$ follows from the pointwise inequality $|\psi_{\mathrm{int}}|\leq C\left||\psi_{\mathrm{int}}|-1\right|$ valid on the support of $1-\chi(|\psi|)$ and
$\|\nabla\psi_{\mathrm{int}}\|_{L^{2}({\mathbf{R}}^{d})}\leq C\|\nabla\psi\|_{L^{2}({\mathbf{R}}^{d})}.$
@@
by means of Minkowski’s inequality. It remains to prove that $(\mathbb{E},d_{\mathbb{E}})$ is a complete metric space. One readily verifies that $d_{\mathbb{E}}$ defines a distance function on $\mathbb{E}({\mathbf{R}}^{d})$. To check that $(\mathbb{E},d_{\mathbb{E}})$ is complete, let $\{\psi_{n}\}_{n}\subset\mathbb{E}$ be a Cauchy sequence with respect to $d_{\mathbb{E}}$. Then, there exists $\psi\in X^{1}+H^{1}$ such that $\psi_{n}\rightarrow\psi$ strongly in $X^{1}+H^{1}$. By lower semi-continuity of norms and (1.10) it follows that $\psi\in\mathbb{E}$. ∎
@@ 2.1. The structure of the energy space depending on the dimension
The structure of the energy space $\mathbb{E}({\mathbf{R}}^{d})$ is sensitive to the dimension $d$. To illustrate this, we recall the following fact. Let $\phi\in\mathcal{D}^{\prime}({\mathbf{R}}^{d})$, if $\nabla\phi\in L^{p}({\mathbf{R}}^{d})$ for some $p<d$, then there exists $c\in{\mathbf{C}}$ such that $\phi-c\in L^{p^{\ast}}({\mathbf{R}}^{d})$, where $p^{\ast}=\frac{dp}{d-p}$, see for instance [41, Theorem 4.5.9]. Hence, if $\psi\in\mathbb{E}({\mathbf{R}}^{3})$, then $\psi$ admits a decomposition $\psi=c+v$ where $c\in{\mathbf{C}}$ with $|c|=1$ and $v\in\dot{H}^{1}({\mathbf{R}}^{3})$, where
@@
denotes the completion of $C_{0}^{\infty}({\mathbf{R}}^{3})$ with respect to the $L^{2}$ norm of the gradient. This observation allows for a equivalent definition of $\mathbb{E}({\mathbf{R}}^{3})$. We equip the space $\mathcal{F}_{c}$ defined in (1.16) with
$\tilde{\delta}(u,v)=\|\nabla u-\nabla v\|_{L^{2}({\mathbf{R}}^{3})}+\||u|^{2}+2\operatorname{Re}(c^{-1}u)-2\operatorname{Re}(c^{-1}v)-|v|^{2}\|_{L^{2}({\mathbf{R}}^{3})}.$
It is straightforward to verify that $\delta$ defines distance function on $\mathcal{F}_{c}$. One has the following characterization given by [26, Proposition 4.1].
@@
In [26], the Proposition is stated for $(\mathbb{E}_{\mathrm{GL}},d_{\mathbb{E}_{\mathrm{GL}}})$. We prove below, see Lemma 2.6, that the two metric spaces can be identified and that the topologies induced by the respective metrics are equivalent.
@@
We observe that the connected components of $\mathbb{E}({\mathbf{R}}^{3})$ are given by $c+\mathcal{F}_{c}({\mathbf{R}}^{3})$ for $c\in{\mathbf{C}}$ with $|c|=1$. The energy space $\mathbb{E}({\mathbf{R}}^{3})$ is an affine space and the far-field behavior is determined by $c$ corresponding to a phase shift. The affine structure of the energy space allows for an alternative approach to solve the Cauchy Problem for $d=3$, as observed in [26, Remark 4.5] for (1.6) and exploited in [47] for cubic-quintic nonlinearities and far-field behavior (1.2).
@@
The $2D$ energy space $\mathbb{E}({\mathbf{R}}^{2})$ lacks an affine structure due to non-trivial oscillations at spatial infinity. Indeed, unbounded phase oscillations at spatial infinity may occur, e.g. $\psi(x)=\mathrm{e}^{\mathrm{i}(2+\log|x|)^{\beta}}$ with $\beta<\frac{1}{2}$ is such that $\psi\in\mathbb{E}({\mathbf{R}}^{2})$, see [26, Remark 4.2]. Moreover, the metric space $(\mathbb{E}({\mathbf{R}}^{2}),d_{\mathbb{E}})$ is not separable. We refer to Remark 2.7 for a detailed discussion and a weakened topology for which $\mathbb{E}({\mathbf{R}}^{2})$ is connected and separable.
@@ 2.2. The Hamiltonian for wave-functions in the energy space
The following inequality turns out to be handy for applications in the sequel. For any $q\in[1,\infty)$ there exists $C_{q}>0$ such that for all $\phi\in L_{\operatorname{loc}}^{1}({\mathbf{R}}^{2})$ with $\mathcal{L}^{2}(\operatorname{supp}(\phi))<+\infty$ and $\nabla\phi\in L^{2}({\mathbf{R}}^{2})$, we have
@@
see for instance [17, Proof of Lemma 2.1]. For $\psi\in\mathbb{E}({\mathbf{R}}^{2})$, applying (2.10) to $\phi=\psi(1-\eta(|\psi|))$ with $\eta$ as defined in (2.3) yields $\psi(1-\eta(|\psi|))\in L^{q}({\mathbf{R}}^{2})$ for any $q\in[1,\infty)$. Indeed, in view of (2.5) it suffices to check that
$\nabla\left(\psi(1-\eta(|\psi|))\right)=(1-\eta(|\psi|))\nabla\psi-\eta^{\prime}(|\psi|)\psi\nabla|\psi|\in L^{2}({\mathbf{R}}^{2})$
since $(1-\eta(|\psi|))\in L^{\infty}({\mathbf{R}}^{2})$, $\psi\eta^{\prime}(|\psi|)\in L^{\infty}({\mathbf{R}}^{2})$ as well as $|\nabla|\psi||\leq|\nabla\psi|$ a.e. on ${\mathbf{R}}^{2}$.
Next, we show that the functional $\mathcal{H}(\psi)$, introduced in (1.10), is bounded for all $\psi\in\mathbb{E}({\mathbf{R}}^{d})$ provided that Assumption 1.1 holds. Specifically, we have $F(1)=0$ and $F^{\prime}(1)=f(1)=0$ which allows one to control $F(|\psi|^{2})\leq C(|\psi|^{2}-1)^{2}$ for $|\psi|^{2}$ sufficiently close to $1$. Furthermore, if one also requires Assumption 1.5 ($f^{\prime}(1)>0$) to be satisfied, then it follows from Taylor expansion that
(2.11) $F(r)\simeq\frac{1}{2}f^{\prime}(1)(r-1)^{2}$
in a small neighborhood of $1$. Hence, there exists $\delta>0$ and $C_{1},C_{2}>0$ such that
provided that $||\psi|^{2}-1|<\delta$. Note that the following Lemma only requires the upper bound on $F$ close to $1$.
@@
Note that we do not require Assumption 1.5 to hold for Lemma 2.5.
@@
In view of (K1) Assumption 1.1, a Taylor expansion of $F$ in a small neighborhood $\mathcal{O}$ of $1$ yields that there exist $C,C^{\prime}>0$ such that
@@
for all $x\in{\mathbf{R}}^{d}$ such that $|\psi|^{2}\in\mathcal{O}$. Let $\delta>0$ be such that $B(1,\delta)\subset\mathcal{O}$ and $\eta_{\delta}(r):=\eta(\frac{r}{\delta})$ with $\eta$ as in (2.3) and $\psi\in\mathbb{E}({\mathbf{R}}^{d})$, then
@@
where we used (K2) Assumption 1.1 in the last inequality. To control the second term, we consider separately the cases $d=2,3$. For $d=3$, Proposition 2.2 yields that there exist $c\in{\mathbf{C}}$ with $|c|=1$ and $v\in\mathcal{F}_{c}({\mathbf{R}}^{3})$ such that $\psi=c+v$ and
@@
where we used (2.5) in the second to last inequality and that $0<\alpha<2$ for $d=3$. For $d=2$, one has that
@@
Next, we identify suitable conditions on $f$ under which the converse inclusion, namely $\{\psi:|\mathcal{H}(\psi)|<+\infty\}\subset\mathbb{E}({\mathbf{R}}^{d})$, holds true. First, we treat the particular case of the Gross-Pitaevskii equation (1.6) for which $\mathcal{H}(\psi)=\mathcal{E}_{\mathrm{GL}}(\psi)$ and so $\mathbb{E}_{\mathrm{GL}}({\mathbf{R}}^{d})=\{\mathcal{H}(\psi)<+\infty\}$, see (1.7) and (1.8) respectively. It has been shown in [26], see also [27], that $(\mathbb{E}_{\mathrm{GL}},d_{\mathbb{E}_{\mathrm{GL}}})$ with
@@
is a complete metric space. It is pointed out in [17, p.13] without proof that $\mathbb{E}=\mathbb{E}_{\mathrm{GL}}$ with equivalent respective metrics. We provide a proof for the sake of completeness.
@@
Lemma 2.6 allows to infer the topological properties of $(\mathbb{E},d_{\mathbb{E}})$ from the results for $(\mathbb{E}_{\mathrm{GL}}({\mathbf{R}}^{d}),d_{\mathbb{E}_{\mathrm{GL}}})$ in [26, 27]. For instance, the functional $\mathcal{E}$ measures the distance to the circle of constants $S^{1}=\{\psi\in\mathbb{E}:\mathcal{E}(\psi)=0\}$ for $d=3$ but not for $d=2$. Indeed, it follows from Lemma 2.6 and [26, Proposition 4.3] that there exists $A>0$ such that for every $\psi\in\mathbb{E}({\mathbf{R}}^{3})$,
@@
If $d=2$, there exists a sequence $\{\psi_{n}\}$ in $\mathbb{E}({\mathbf{R}}^{2})$ such that $\mathcal{E}(\psi_{n})\rightarrow 0$ but $d_{\mathbb{E}}(\psi_{n},S^{1})\geq c_{0}>0$. Note that the complete metric space $(\mathbb{E}_{\mathrm{GL}}({\mathbf{R}}^{2}),d_{\mathbb{E}_{\mathrm{GL}}})$ lacks an affine structure and to be separable. In [27] a detailed characterisation of $\mathbb{E}_{\mathrm{GL}}({\mathbf{R}}^{d})$ including a manifold structure for $\mathbb{E}_{\mathrm{GL}}({\mathbf{R}}^{d})$ is provided. The connected components are characterised by [27, Theorem 1.8] and [27, Proposition 1.10]. We introduce a (strictly) weaker topology [27, p. 140] induced by the metric